Consider the 3-dimensional vector field F defined by F(x,y,z)=(2xyz,x²z+2yz²,x²y+2y²z+e^z).


1.write down the Jacobian matrix jf(x,y,z).


2.determine divF (x,y,z).


3.determine curl F (x,y,z).


4.does F have a potential function? Give reasons for your answer, referring to the relevant definitions and theorems in the study guide.


5.find a potential function of F .

1
Expert's answer
2022-06-21T08:09:02-0400

1.


"F(x,y,z)=(2xyz,x^2z+2yz^2,x^2y+2y^2z+e^z)"


"\dfrac{\partial F_x}{\partial x}=2yz, \dfrac{\partial F_y}{\partial x}=2xz, \dfrac{\partial F_z}{\partial x}=2xy"


"\dfrac{\partial F_x}{\partial y}=2xz, \dfrac{\partial F_y}{\partial y}=2z^2, \dfrac{\partial F_z}{\partial y}=x^2+4yz"

"\dfrac{\partial F_x}{\partial z}=2xy, \dfrac{\partial F_y}{\partial z}=x^2+4yz, \dfrac{\partial F_z}{\partial z}=2y^2+e^z"

Jacobian Matrix


"J(F(x,y,z))=\begin{vmatrix}
 2yz & 2xz & 2xy \\
 2xz & 2z^2 & x^2+4yz \\
 2xy & x^2+4yz & 2y^2+e^z \\
\end{vmatrix}"


"=2yz\begin{vmatrix}
 2z^2 & x^2+4yz \\
 x^2+4yz & 2y^2+e^z
\end{vmatrix}"

"-2xz\begin{vmatrix}
 2xz & x^2+4yz \\
 2xy & 2y^2+e^z
\end{vmatrix}"

"+2xy\begin{vmatrix}
 2xz & 2z^2 \\
 2xy & x^2+4yz
\end{vmatrix}"

"=8y^3z^3+4yz^3e^z-2x^4yz-16x^2y^2z^2-32y^3z^3"

"-8x^2y^2z^2-4x^2z^2e^z+4x^4yz+16x^2y^2z^2"

"+4x^4yz+16x^2y^2z^2-8x^2y^2z^2"

"=6x^4yz-24y^3z^3-4x^2z^2e^z+4yz^3e^z"

2.


"\nabla=\langle\dfrac{\partial }{\partial x},\dfrac{\partial }{\partial y},\dfrac{\partial }{\partial z}\rangle"


"divF=\nabla F=2yz+2z^2+2y^2+e^z"

3.


"curl F(x,y,z)=\begin{vmatrix}
 \vec{i} & \vec{j} & \vec{k} \\
 \dfrac{\partial }{\partial x} & \dfrac{\partial }{\partial y} & \dfrac{\partial }{\partial z} \\
2xyz & x^2z+2yz^2 & x^2y+2y^2z+e^z

\end{vmatrix}"

"=\vec{i}(x^2+4yz-x^2-4yz)"

"-\vec{j}(2xy-2xy)"


"+\vec{k}(2xz-2xz)=\vec{0}"


4. Let "F(x,y,z)"  be a vector field in space on a simply connected domain. If  "curl F(x,y,z)=0," then "F" is conservative.

There exists a function "f" such that "F=\nabla f."In this situation "f" is called

a potential function for "F."


5.


"f_x=2xyz=>f=x^2yz+g(y,z)"

"f_y=x^2z+g_y=x^2z+2yz^2"




"g_y=2yz^2"

"g=y^2z^2+h(z)"

"f=x^2yz+y^2z^2+h(z)"


"f_z=x^2y+2y^2z+h'(z)=x^2y+2y^2z+e^z"

"h'(z)=e^z=>h(z)=e^z+C"

A potential function of F is


"f=x^2yz+y^2z^2+e^z+C"